# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 4.3. The Fefferman space for twistor CR manifolds
We will compute the pull-back $\varpi^{*}g^{\mathrm{F}}$ of the Fefferman metric, which we also denote by $g^{\mathrm{F}}$ and call the Fefferman metric on $\widehat{\mathcal{C}}$. We set
$\hat{\xi}:=\varpi^{*}\xi=\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2},\quad\hat{\varphi}:=\varpi^{*}\varphi,\quad\hat{\sigma}:=\varpi^{*}\sigma,$
where $\xi,\varphi,\sigma$ are the differential forms defined in §2.2.
@@
$d\hat{\xi}$ $=4i\hat{\sigma}\wedge\hat{\xi},$
$\hat{\sigma}\wedge d\hat{\varphi}\wedge\overline{\hat{\varphi}}$ $=(\mathrm{Tr}\,d\hat{\sigma})i\hat{\sigma}\wedge\theta\wedge\hat{\varphi}\wedge\overline{\hat{\varphi}}.$
@@
$d\hat{\xi}=2i\tilde{\theta}^{3}\wedge\hat{\xi},\quad\tilde{\theta}^{3}\wedge d\hat{\varphi}\wedge\overline{\hat{\varphi}}=0.$
@@
$d\tilde{\theta}^{3}$ $=2iP(\zeta,\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+i\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}}+\cdots$
$=2iP(\zeta,\overline{\zeta})\theta^{1}\wedge\theta^{\overline{1}}+i\theta^{2}\wedge\theta^{\overline{2}}+\cdots,$
we have ${\rm Tr}\,d\tilde{\theta}^{3}=0$. Thus, we obtain $\hat{\sigma}=\frac{1}{2}\tilde{\theta}^{3}$. ∎
Since $g^{\mathrm{F}}=2h_{\alpha\overline{\beta}}\theta^{\alpha}\cdot\theta^{\overline{\beta}}+4\theta\cdot\hat{\sigma}=2h_{\alpha\overline{\beta}}\tilde{\theta}^{\alpha}\cdot\tilde{\theta}^{\overline{\beta}}+4\theta\cdot\hat{\sigma}$, we obtain the following theorem:
@@
$g^{\mathrm{F}}=2(\tilde{\theta}^{1}\cdot\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\cdot\tilde{\theta}^{\overline{1}}+\theta\cdot\tilde{\theta}^{3})$
@@
The non-zero components of $g^{\mathrm{F}}$ in the adapted coframe $(\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3})$ are
(4.3) $g^{\mathrm{F}}_{1\overline{2}}=g^{\mathrm{F}}_{\overline{2}1}=g^{\mathrm{F}}_{2\overline{1}}=g^{\mathrm{F}}_{\overline{1}2}=g^{\mathrm{F}}_{03}=g^{\mathrm{F}}_{30}=1.$
4.4. The connection form of $g^{\mathrm{F}}$
The connection 1-forms $\omega_{a}{}^{b}$ of the Fefferman metric are computed as follows:
@@ 4.4. The connection form of $g^{\mathrm{F}}$
In the coframe $\{\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\}$, the connection forms of the Levi-Civita connection of $g^{\rm F}$ are given by
$\omega_{0}{}^{0}$ $=0,$ $\omega_{1}{}^{0}$ $=\frac{i}{2}\tilde{\theta}^{\overline{2}},$ $\omega_{2}{}^{0}$ $=\frac{i}{2}\tilde{\theta}^{\overline{1}},$ $\omega_{3}{}^{0}$ $=0,$
$\omega_{0}{}^{1}$ $=\frac{i}{2}\tilde{\theta}^{2},$ $\omega_{1}{}^{1}$ $=-\frac{i}{2}\tilde{\theta}^{3},$ $\omega_{2}{}^{1}$ $=-\frac{i}{2}\theta,$ $\omega_{\overline{1}}{}^{1}$ $=0,$ $\omega_{\overline{2}}{}^{1}$ $=0,$ $\omega_{3}{}^{1}$ $=\frac{i}{2}\tilde{\theta}^{1},$
$\omega_{0}{}^{2}$ $=iP(\zeta,\eta)\theta+iP(\zeta,\overline{\zeta})\tilde{\theta}^{1}+iP(\zeta,\zeta)\tilde{\theta}^{\overline{1}},$
$\omega_{1}{}^{2}$ $=-iP(\eta,\eta)\theta-iP(\overline{\zeta},\eta)\tilde{\theta}^{1}-iP(\zeta,\eta)\tilde{\theta}^{\overline{1}},$
$\omega_{2}{}^{2}$ $=-\frac{i}{2}\tilde{\theta}^{3},\quad\omega_{\overline{1}}{}^{2}=0,\quad\omega_{\overline{2}}{}^{2}=0,\quad\omega_{3}{}^{2}=\frac{i}{2}\tilde{\theta}^{2},$
$\omega_{0}{}^{3}$ $=0,\quad\omega_{1}{}^{3}=iP(\overline{\zeta},\eta)\theta+iP(\overline{\zeta},\overline{\zeta})\tilde{\theta}^{1}+iP(\zeta,\overline{\zeta})\tilde{\theta}^{\overline{1}},\quad\omega_{2}{}^{3}=\frac{i}{2}\tilde{\theta}^{\overline{2}},\quad\omega_{3}{}^{3}=0$
@@
$g^{\mathrm{F}}_{bc}\omega_{a}{}^{c}+g^{\mathrm{F}}_{ac}\omega_{b}{}^{c}=dg^{\mathrm{F}}_{ab},\quad d\tilde{\theta}^{a}=\tilde{\theta}^{b}\wedge\omega_{b}{}^{a}.$
@@
(4.4) $\omega_{0}{}^{3}$ $=0,$ $\omega_{0}{}^{\overline{2}}+\omega_{1}{}^{3}$ $=0,$ $\omega_{0}{}^{\overline{1}}+\omega_{2}{}^{3}$ $=0,$ $\omega_{0}{}^{0}+\omega_{3}{}^{3}$ $=0,$
$\omega_{1}{}^{\overline{2}}$ $=0,$ $\omega_{1}{}^{\overline{1}}+\omega_{2}{}^{\overline{2}}$ $=0,$ $\omega_{1}{}^{2}+\omega_{\overline{1}}{}^{\overline{2}}$ $=0,$ $\omega_{1}{}^{1}+\omega_{\overline{2}}{}^{\overline{2}}$ $=0,$
$\omega_{1}{}^{0}+\omega_{3}{}^{\overline{2}}$ $=0,$ $\omega_{2}{}^{\overline{1}}$ $=0,$ $\omega_{2}{}^{1}+\omega_{\overline{2}}{}^{\overline{1}}$ $=0,$ $\omega_{2}{}^{0}+\omega_{3}{}^{\overline{1}}$ $=0,$
$\omega_{3}{}^{0}$ $=0$
@@
$d\tilde{\theta}^{a}=\theta\wedge\omega_{0}{}^{a}+\tilde{\theta}^{1}\wedge\omega_{1}{}^{a}+\tilde{\theta}^{2}\wedge\omega_{2}{}^{a}+\tilde{\theta}^{\overline{1}}\wedge\omega_{\overline{1}}{}^{a}+\tilde{\theta}^{\overline{2}}\wedge\omega_{\overline{2}}{}^{a}+\tilde{\theta}^{3}\wedge\omega_{3}{}^{a}$
@@
4.5. The curvature form of $g^{\mathrm{F}}$
@@
$\Omega_{a}{}^{b}=d\omega_{a}{}^{b}-\omega_{a}{}^{c}\wedge\omega_{c}{}^{b}=\frac{1}{2}R^{\mathrm{F}}_{cd}{}^{b}{}_{a}\tilde{\theta}^{c}\wedge\tilde{\theta}^{d}$
by using Proposition 4.6 and the equations (4.1), (4.2). Since $\Omega_{a}{}^{b}$ satisfies the same symmetries as in (4.4), we present only
$\Omega_{0}{}^{0},\ \Omega_{1}{}^{0},\ \Omega_{2}{}^{0},\ \Omega_{0}{}^{1},\ \Omega_{2}{}^{1},\ \Omega_{0}{}^{2},\ \Omega_{2}{}^{2},\ \Omega_{1}{}^{2},\ \Omega_{2}{}^{\overline{2}}.$
@@
$\Omega_{0}{}^{0}$ $=\frac{1}{2}P(\overline{\zeta},\eta)\theta\wedge\tilde{\theta}^{1}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}},$
$\Omega_{1}{}^{0}$ $=\frac{1}{2}P(\overline{\zeta},\overline{\zeta})\theta\wedge\tilde{\theta}^{1}+\frac{1}{2}P(\zeta,\overline{\zeta})\theta\wedge\tilde{\theta}^{\overline{1}}-\frac{1}{4}\tilde{\theta}^{\overline{2}}\wedge\tilde{\theta}^{3},$
@@
$\Omega_{0}{}^{2}$ $=(*C)(\zeta,\overline{\zeta})\theta\wedge\tilde{\theta}^{1}-\frac{1}{2}P(\eta,\eta)\theta\wedge\tilde{\theta}^{2}-(*C)(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{3}$
$\quad-\frac{1}{2}P(\overline{\zeta},\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{2}+(*C)(\zeta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{3}+\frac{1}{2}P(\zeta,\eta)\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}$
$\quad+\frac{1}{2}P(\zeta,\zeta)\tilde{\theta}^{\overline{1}}\wedge\tilde{\theta}^{3},$
$\Omega_{2}{}^{2}$ $=\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{4}\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}},$
$\Omega_{1}{}^{2}$ $=-(*C)(\overline{\zeta},\eta)\theta\wedge\tilde{\theta}^{1}-\frac{1}{2}P(\overline{\zeta},\eta)\theta\wedge\tilde{\theta}^{2}+(*C)(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{2}}$
$\quad-\frac{1}{2}P(\overline{\zeta},\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{2}-(*C)(\eta,\eta)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{2}}+\frac{1}{2}P(\zeta,\overline{\zeta})\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{1}}$
$\quad+\frac{1}{2}P(\zeta,\zeta)\tilde{\theta}^{\overline{1}}\wedge\tilde{\theta}^{\overline{2}},$
$\Omega_{2}{}^{\overline{2}}$ $=\frac{1}{2}P(\overline{\zeta},\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\overline{\zeta},\overline{\zeta})\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}.$
@@
$R^{\mathrm{F}}_{010\overline{2}}$ $=\frac{1}{2}P(\eta,\eta),$ $R^{\mathrm{F}}_{0\overline{2}1\overline{1}}$ $=-\frac{1}{2}P(\zeta,\eta),$ $R^{\mathrm{F}}_{0\overline{2}23}$ $=-\frac{1}{4},$
$R^{\mathrm{F}}_{010\overline{1}}$ $=-(*C)(\zeta,\overline{\zeta}),$ $R^{\mathrm{F}}_{0\overline{1}0\overline{1}}$ $=(*C)(\zeta,\zeta),$ $R^{\mathrm{F}}_{0\overline{1}03}$ $=-\frac{1}{2}P(\zeta,\eta),$
$R^{\mathrm{F}}_{0\overline{1}12}$ $=\frac{1}{2}P(\overline{\zeta},\eta),$ $R^{\mathrm{F}}_{0\overline{1}2\overline{1}}$ $=-\frac{1}{2}P(\zeta,\eta),$ $R^{\mathrm{F}}_{0\overline{1}1\overline{1}}$ $=-(*C)(\zeta,\eta),$
$R^{\mathrm{F}}_{0\overline{1}13}$ $=-\frac{1}{2}P(\zeta,\overline{\zeta}),$ $R^{\mathrm{F}}_{0\overline{1}\overline{1}3}$ $=-\frac{1}{2}P(\zeta,\zeta),$ $R^{\mathrm{F}}_{121\overline{1}}$ $=\frac{1}{2}P(\overline{\zeta},\overline{\zeta}),$
$R^{\mathrm{F}}_{2\overline{1}2\overline{2}}$ $=-\frac{1}{4},$ $R^{\mathrm{F}}_{1\overline{1}2\overline{1}}$ $=-\frac{1}{2}P(\zeta,\overline{\zeta}),$ $R^{\mathrm{F}}_{23\overline{1}3}$ $=\frac{1}{4},$
@@
$R^{\mathrm{F}}_{00}$ $=2P(\eta,\eta),$ $R^{\mathrm{F}}_{01}$ $=2P(\overline{\zeta},\eta),$ $R^{\mathrm{F}}_{2\overline{2}}$ $=1,$
$R^{\mathrm{F}}_{11}$ $=2P(\overline{\zeta},\overline{\zeta}),$ $R^{\mathrm{F}}_{1\overline{1}}$ $=2P(\zeta,\overline{\zeta}),$ $R^{\mathrm{F}}_{33}$ $=1.$
It follows that the scalar curvature vanishes:
$R^{\mathrm{F}}=2(R^{\mathrm{F}}_{1\overline{2}}+R^{\mathrm{F}}_{2\overline{1}}+R^{\mathrm{F}}_{03})=0.$
@@
$W^{\mathrm{F}}_{abcd}=R^{\mathrm{F}}_{abcd}-\frac{1}{4}(R^{\mathrm{F}}_{ac}g^{\mathrm{F}}_{bd}-R^{\mathrm{F}}_{bc}g^{\mathrm{F}}_{ad}+R^{\mathrm{F}}_{bd}g^{\mathrm{F}}_{ac}-R^{\mathrm{F}}_{ad}g^{\mathrm{F}}_{bc})$
@@
$W^{\mathrm{F}}_{010\overline{1}}$ $=-(*C)(\zeta,\overline{\zeta}),$ $W^{\mathrm{F}}_{0\overline{1}0\overline{1}}$ $=(*C)(\zeta,\zeta),$
$W^{\mathrm{F}}_{0\overline{1}1\overline{1}}$ $=-(*C)(\zeta,\eta),$ $W^{\mathrm{F}}_{1\overline{1}1\overline{1}}$ $=(*C)(\eta,\eta)$
and the components which can be obtained from these by the symmetry and reality of the Weyl tensor. Since $(\zeta,\overline{\zeta},\eta)$ gives a basis of $\mathbb{C}T\Sigma$ at the base point, we obtain the following theorem:
Theorem 4.8 ([19, Theorem 6.7]).
The Fefferman space $(\widehat{\mathcal{C}},g^{\mathrm{F}})$ is conformally flat if and only if $(\Sigma,[g])$ is conformally flat.
@@
4.6. Null geodesics of $g^{\mathrm{F}}$
To examine the projections of chains and null chains to $\Sigma$, we compute the null geodesic equation for the Fefferman metric.
We fix a representative metric $g\in[g]$, and let $\gamma(t)=(x^{i}(t),\zeta_{i}(t))$ be a curve on $\widehat{\mathcal{C}}$ such that the projection $x(t)=(x^{i}(t))$ is a regular curve on $\Sigma$. In the adapted coframe $\{\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\}$, the components $\dot{\gamma}^{a}:=\tilde{\theta}^{a}(\dot{\gamma})$ are given by

### v4

@@ 4.3 The Fefferman space for twistor CR manifolds
We will compute the pullback $\varpi^{*}g^{\mathrm{F}}$ of the Fefferman metric, which we also denote by $g^{\mathrm{F}}$ and call the Fefferman metric on $\widehat{\mathcal{C}}$. We set
$\hat{\xi}:=\varpi^{*}\xi=\theta\wedge\tilde{\theta}^{1}\wedge\tilde{\theta}^{2},\qquad\hat{\varphi}:=\varpi^{*}\varphi,\qquad\hat{\sigma}:=\varpi^{*}\sigma,$
where $\xi$, $\varphi$, $\sigma$ are the differential forms defined in Section 2.2.
@@
${\rm d}\hat{\xi}=4{\rm i}\hat{\sigma}\wedge\hat{\xi}$,
$\hat{\sigma}\wedge{\rm d}\hat{\varphi}\wedge\overline{\hat{\varphi}}=(\operatorname{Tr}{\rm d}\hat{\sigma}){\rm i}\hat{\sigma}\wedge\theta\wedge\hat{\varphi}\wedge\overline{\hat{\varphi}}$.
@@
${\rm d}\hat{\xi}=2{\rm i}\tilde{\theta}^{3}\wedge\hat{\xi},\qquad\tilde{\theta}^{3}\wedge{\rm d}\hat{\varphi}\wedge\overline{\hat{\varphi}}=0.$
@@
${\rm d}\tilde{\theta}^{3}=2{\rm i}P\bigl(\zeta,\overline{\zeta}\bigr)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+{\rm i}\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}}+\cdots=2{\rm i}P\bigl(\zeta,\overline{\zeta}\bigr)\theta^{1}\wedge\theta^{\overline{1}}+{\rm i}\theta^{2}\wedge\theta^{\overline{2}}+\cdots,$
we have $\operatorname{Tr}{\rm d}\tilde{\theta}^{3}=0$. Thus, we obtain $\hat{\sigma}=\frac{1}{2}\tilde{\theta}^{3}$. ∎
Since $g^{\mathrm{F}}=2h_{\alpha\overline{\beta}}\theta^{\alpha}\cdot\theta^{\overline{\beta}}+4\theta\cdot\hat{\sigma}=2h_{\alpha\overline{\beta}}\tilde{\theta}^{\alpha}\cdot\tilde{\theta}^{\overline{\beta}}+4\theta\cdot\hat{\sigma}$, we obtain the following theorem.
@@
$g^{\mathrm{F}}=2\bigl(\tilde{\theta}^{1}\cdot\tilde{\theta}^{\overline{2}}+\tilde{\theta}^{2}\cdot\tilde{\theta}^{\overline{1}}+\theta\cdot\tilde{\theta}^{3}\bigr)$
@@
The non-zero components of $g^{\mathrm{F}}$ in the adapted coframe $\bigl(\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr)$ are
$g^{\mathrm{F}}_{1\overline{2}}=g^{\mathrm{F}}_{\overline{2}1}=g^{\mathrm{F}}_{2\overline{1}}=g^{\mathrm{F}}_{\overline{1}2}=g^{\mathrm{F}}_{03}=g^{\mathrm{F}}_{30}=1.$ (4.3)
4.4 The connection form of $\boldsymbol{g^{\mathrm{F}}}$
The connection 1-forms $\omega_{a}{}^{b}$ of the Fefferman metric are computed as follows.
@@ 4.4 The connection form of $\boldsymbol{g^{\mathrm{F}}}$
In the coframe $\bigl(\tilde{\theta}^{0}=\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr)$, the connection forms of the Levi-Civita connection of $g^{\rm F}$ are given by
[…]
$\omega_{0}{}^{2}={\rm i}P(\zeta,\eta)\theta+{\rm i}P\bigl(\zeta,\overline{\zeta}\bigr)\tilde{\theta}^{1}+{\rm i}P(\zeta,\zeta)\tilde{\theta}^{\overline{1}},$
$\omega_{1}{}^{2}=-{\rm i}P(\eta,\eta)\theta-{\rm i}P\bigl(\overline{\zeta},\eta\bigr)\tilde{\theta}^{1}-{\rm i}P(\zeta,\eta)\tilde{\theta}^{\overline{1}},$
$\omega_{2}{}^{2}=-\frac{{\rm i}}{2}\tilde{\theta}^{3},\qquad\omega_{\overline{1}}{}^{2}=0,\qquad\omega_{\overline{2}}{}^{2}=0,\qquad\omega_{3}{}^{2}=\frac{{\rm i}}{2}\tilde{\theta}^{2},$
$\omega_{0}{}^{3}=0,\qquad\omega_{1}{}^{3}={\rm i}P\bigl(\overline{\zeta},\eta\bigr)\theta+{\rm i}P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\tilde{\theta}^{1}+{\rm i}P\bigl(\zeta,\overline{\zeta}\bigr)\tilde{\theta}^{\overline{1}},\qquad\omega_{2}{}^{3}=\frac{{\rm i}}{2}\tilde{\theta}^{\overline{2}},\qquad\omega_{3}{}^{3}=0$
@@
$g^{\mathrm{F}}_{bc}\omega_{a}{}^{c}+g^{\mathrm{F}}_{ac}\omega_{b}{}^{c}=dg^{\mathrm{F}}_{ab},\qquad{\rm d}\tilde{\theta}^{a}=\tilde{\theta}^{b}\wedge\omega_{b}{}^{a}$.
@@
$\omega_{0}{}^{3}=0,\qquad$ $\omega_{0}{}^{\overline{2}}+\omega_{1}{}^{3}=0,\qquad$ $\omega_{0}{}^{\overline{1}}+\omega_{2}{}^{3}=0,\qquad$ $\omega_{0}{}^{0}+\omega_{3}{}^{3}=0,$
$\omega_{1}{}^{\overline{2}}=0,\qquad$ $\omega_{1}{}^{\overline{1}}+\omega_{2}{}^{\overline{2}}=0,\qquad$ $\omega_{1}{}^{2}+\omega_{\overline{1}}{}^{\overline{2}}=0,\qquad$ $\omega_{1}{}^{1}+\omega_{\overline{2}}{}^{\overline{2}}=0,$
$\omega_{1}{}^{0}+\omega_{3}{}^{\overline{2}}=0,\qquad$ $\omega_{2}{}^{\overline{1}}=0,\qquad$ $\omega_{2}{}^{1}+\omega_{\overline{2}}{}^{\overline{1}}=0,\qquad$ $\omega_{2}{}^{0}+\omega_{3}{}^{\overline{1}}=0,$
$\omega_{3}{}^{0}=0$ (4.4)
@@
${\rm d}\tilde{\theta}^{a}=\theta\wedge\omega_{0}{}^{a}+\tilde{\theta}^{1}\wedge\omega_{1}{}^{a}+\tilde{\theta}^{2}\wedge\omega_{2}{}^{a}+\tilde{\theta}^{\overline{1}}\wedge\omega_{\overline{1}}{}^{a}+\tilde{\theta}^{\overline{2}}\wedge\omega_{\overline{2}}{}^{a}+\tilde{\theta}^{3}\wedge\omega_{3}{}^{a}$
@@
4.5 The curvature form of $\boldsymbol{g^{\mathrm{F}}}$
@@
$\Omega_{a}{}^{b}={\rm d}\omega_{a}{}^{b}-\omega_{a}{}^{c}\wedge\omega_{c}{}^{b}=\frac{1}{2}R^{\mathrm{F}}_{cd}{}^{b}{}_{a}\tilde{\theta}^{c}\wedge\tilde{\theta}^{d}$
by using Proposition 4.6 and the equations (4.1) and (4.2). Since $\Omega_{a}{}^{b}$ satisfies the same symmetries as in (4.4), we present only $\Omega_{0}{}^{0}$, $\Omega_{1}{}^{0}$, $\Omega_{2}{}^{0}$, $\Omega_{0}{}^{1}$, $\Omega_{2}{}^{1}$, $\Omega_{0}{}^{2}$, $\Omega_{2}{}^{2}$, $\Omega_{1}{}^{2}$, $\Omega_{2}{}^{\overline{2}}$.
@@
$\Omega_{0}{}^{0}=\frac{1}{2}P\bigl(\overline{\zeta},\eta\bigr)\theta\wedge\tilde{\theta}^{1}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}},$
$\Omega_{1}{}^{0}=\frac{1}{2}P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\theta\wedge\tilde{\theta}^{1}+\frac{1}{2}P\bigl(\zeta,\overline{\zeta}\bigr)\theta\wedge\tilde{\theta}^{\overline{1}}-\frac{1}{4}\tilde{\theta}^{\overline{2}}\wedge\tilde{\theta}^{3},$
@@
$\Omega_{0}{}^{2}=(*C)\bigl(\zeta,\overline{\zeta}\bigr)\theta\wedge\tilde{\theta}^{1}-\frac{1}{2}P(\eta,\eta)\theta\wedge\tilde{\theta}^{2}-(*C)(\zeta,\zeta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{3}$
[…]
$\Omega_{2}{}^{2}=\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P\bigl(\zeta,\overline{\zeta}\bigr)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{4}\tilde{\theta}^{2}\wedge\tilde{\theta}^{\overline{2}},$
$\Omega_{1}{}^{2}=-(*C)\bigl(\overline{\zeta},\eta\bigr)\theta\wedge\tilde{\theta}^{1}-\frac{1}{2}P\bigl(\overline{\zeta},\eta\bigr)\theta\wedge\tilde{\theta}^{2}+(*C)(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P(\zeta,\eta)\theta\wedge\tilde{\theta}^{\overline{2}}$
[…]
$\Omega_{2}{}^{\overline{2}}=\frac{1}{2}P\bigl(\overline{\zeta},\eta\bigr)\theta\wedge\tilde{\theta}^{\overline{1}}+\frac{1}{2}P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\tilde{\theta}^{1}\wedge\tilde{\theta}^{\overline{1}}.$
@@
$R^{\mathrm{F}}_{010\overline{2}}=\frac{1}{2}P(\eta,\eta),\qquad$ $R^{\mathrm{F}}_{0\overline{2}1\overline{1}}=-\frac{1}{2}P(\zeta,\eta),\qquad$ $R^{\mathrm{F}}_{0\overline{2}23}=-\frac{1}{4},$
$R^{\mathrm{F}}_{010\overline{1}}=-(*C)\bigl(\zeta,\overline{\zeta}\bigr),\qquad$ $R^{\mathrm{F}}_{0\overline{1}0\overline{1}}=(*C)(\zeta,\zeta),\qquad$ $R^{\mathrm{F}}_{0\overline{1}03}=-\frac{1}{2}P(\zeta,\eta),$
$R^{\mathrm{F}}_{0\overline{1}12}=\frac{1}{2}P\bigl(\overline{\zeta},\eta\bigr),\qquad$ $R^{\mathrm{F}}_{0\overline{1}2\overline{1}}=-\frac{1}{2}P(\zeta,\eta),\qquad$ $R^{\mathrm{F}}_{0\overline{1}1\overline{1}}=-(*C)(\zeta,\eta),$
$R^{\mathrm{F}}_{0\overline{1}13}=-\frac{1}{2}P\bigl(\zeta,\overline{\zeta}\bigr),\qquad$ $R^{\mathrm{F}}_{0\overline{1}\overline{1}3}=-\frac{1}{2}P(\zeta,\zeta),\qquad$ $R^{\mathrm{F}}_{121\overline{1}}=\frac{1}{2}P\bigl(\overline{\zeta},\overline{\zeta}\bigr),$
$R^{\mathrm{F}}_{2\overline{1}2\overline{2}}=-\frac{1}{4},\qquad$ $R^{\mathrm{F}}_{1\overline{1}2\overline{1}}=-\frac{1}{2}P\bigl(\zeta,\overline{\zeta}\bigr),\qquad$ $R^{\mathrm{F}}_{23\overline{1}3}=\frac{1}{4},$
@@
$R^{\mathrm{F}}_{00}=2P(\eta,\eta),\qquad$ $R^{\mathrm{F}}_{01}=2P\bigl(\overline{\zeta},\eta\bigr),\qquad$ $R^{\mathrm{F}}_{2\overline{2}}=1,$
$R^{\mathrm{F}}_{11}=2P\bigl(\overline{\zeta},\overline{\zeta}\bigr),\qquad$ $R^{\mathrm{F}}_{1\overline{1}}=2P\bigl(\zeta,\overline{\zeta}\bigr),\qquad$ $R^{\mathrm{F}}_{33}=1.$
It follows that the scalar curvature vanishes,
$R^{\mathrm{F}}=2\bigl(R^{\mathrm{F}}_{1\overline{2}}+R^{\mathrm{F}}_{2\overline{1}}+R^{\mathrm{F}}_{03}\bigr)=0.$
@@
$W^{\mathrm{F}}_{abcd}=R^{\mathrm{F}}_{abcd}-\frac{1}{4}\bigl(R^{\mathrm{F}}_{ac}g^{\mathrm{F}}_{bd}-R^{\mathrm{F}}_{bc}g^{\mathrm{F}}_{ad}+R^{\mathrm{F}}_{bd}g^{\mathrm{F}}_{ac}-R^{\mathrm{F}}_{ad}g^{\mathrm{F}}_{bc}\bigr)$
@@
$W^{\mathrm{F}}_{010\overline{1}}=-(*C)\bigl(\zeta,\overline{\zeta}\bigr),\qquad$ $W^{\mathrm{F}}_{0\overline{1}0\overline{1}}=(*C)(\zeta,\zeta),$
$W^{\mathrm{F}}_{0\overline{1}1\overline{1}}=-(*C)(\zeta,\eta),\qquad$ $W^{\mathrm{F}}_{1\overline{1}1\overline{1}}=(*C)(\eta,\eta)$
and the components which can be obtained from these by the symmetry and reality of the Weyl tensor. Since $\bigl(\zeta,\overline{\zeta},\eta\bigr)$ gives a basis of $\mathbb{C}T\Sigma$ at the base point, we obtain the following theorem.
Theorem 4.8 ([23, Theorem 6.7] and [25]).
The Fefferman space $\bigl(\widehat{\mathcal{C}},g^{\mathrm{F}}\bigr)$ is conformally flat if and only if $(\Sigma,[g])$ is conformally flat.
@@
4.6 Null geodesics of $\boldsymbol{g^{\mathrm{F}}}$
To examine the projections of chains and null chains to $\Sigma$, we compute the null geodesic equation for the Fefferman metric. We fix a representative metric $g\in[g]$, and let $\gamma(t)=\bigl(x^{i}(t),\zeta_{i}(t)\bigr)$ be a curve on $\widehat{\mathcal{C}}$ such that the projection $x(t)\!=\!\bigl(x^{i}(t)\bigr)$ is a regular curve on $\Sigma$. In the adapted coframe $\smash{\bigl(\tilde{\theta}^{0}\!=\!\theta,\tilde{\theta}^{1},\tilde{\theta}^{2},\tilde{\theta}^{\overline{1}},\tilde{\theta}^{\overline{2}},\tilde{\theta}^{3}\bigr)}$, the components $\dot{\gamma}^{a}:=\tilde{\theta}^{a}(\dot{\gamma})$ are given by